OpenStudy (anonymous):

Multiply. State any excluded values. 5/2a * 3/a^2 A. 15/2a^3 ; where a = 0 B. 15/2a^2 C. 1/a^2 ; where a = 0

OpenStudy (anonymous):

\[\frac{5}{2a}\times \frac{3}{a^2}=\frac{5\times 3}{2aright 6 units (from -3 to 3) down seven units (from 1 to -6) slope is \(-\frac{7}{6}\)\times a^2}=\frac{15}{2a^3}\] and yes, \(x\neq 0\)